OpenStudy (anonymous):

A line passes through (2, –1) and (8, 4). a. Choose the correct equation for the line in point-slope form. b. What is the equation rewritten in standard form?

OpenStudy (anonymous):

use the formula\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

this will give you the slope

OpenStudy (anonymous):

and then (y-y1) = slope*(x-x1) where slope is as said by @sjerman1

OpenStudy (anonymous):

good job @SandeepReddy that is correct @_kaaroo you use the formula that @SandeepReddy gave after you find out the slope

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

y-1=5/6 (x-2); -5x + 6y = 16 @sjerman1

OpenStudy (anonymous):

well you need to put it into point slope form which is y=mx+b but good job so far